OpenStudy (anonymous):

The volume of a rectangular prism is given by the formula V = lwh, where l is the length of the prism, w is the width, and h is the height. Suppose a box in the shape of a rectangular prism has length (2a + 11), width (5a – 12), and height (a + 6). Which expression represents the volume of the box?

OpenStudy (anonymous):

v=(2a+11)(5a-12)(a+6) v=(10a^2-24a+55a-132)(a+6) v=(10a^3+60a^2-24a^2-144a+55a^2+330a-132a-792 v=10a^3+60a^2-24a^2+55a^2-144a+330a-132a-792 v=10^3+91a^2+54a-792